princeevee:

i need some help. find the percentage decrease from 10 1/2 to 8 2/5

princeevee:

@Shadow

princeevee:

@Eiwoh2

theghost22:

Are you still here?

princeevee:

yes

theghost22:

Ok so I am pretty sure that I have figured it out

theghost22:

Give me a moment while I explain

theghost22:

So if you take both fractions and you convert them into a percentage what do you get?

theghost22:

The first one is going to be 1050%

princeevee:

350 and 437

theghost22:

Is rthat what you get when you take them both and put them into a fraction?

theghost22:

Percentage*

princeevee:

mhm

princeevee:

oh wait i need to add them?

theghost22:

It should be 1050% and 840%

princeevee:

oh

theghost22:

You take the fractions and put them into an percentage and then divide the percentage together

theghost22:

So take 1050% and divide it by 840%

princeevee:

1.25

theghost22:

Your answer should be 250%

theghost22:

@Shadow confirm my answer please?

theghost22:

That should be your answer but I just wanted to confirm it

jhonyy9:

my idea 10 1/2 = 21/2 8 2/5 = 42/5 21/2 mean 100% 42/5 mean x% cross multiplie and get x = [(42/5)*100]/(21/2) = (42*20)/(21/2) = (840*2)/21 = =1680/21 = 80% so the percentage decrease from 10 1/2 to 8 2/5 is 100% - 80% = 20% hope this is understandably